Barton Industries estimates its cost of common equity by using three approaches: the CAPM, the bond-yield-plus-risk-premium approach, and the DCF model. Barton expects next year's annual dividend, D1, to be $2.50 and it expects dividends to grow at a constant rate gL = 3.7%. The firm's current common stock price, P0, is $22.00. The current risk-free rate, rRF, = 4.7%; the market risk premium, RPM, = 6%, and the firm's stock has a current beta, b, = 1.2. Assume that the firm's cost of debt, rd, is 9.48%. The firm uses a 4% risk premium when arriving at a ballpark estimate of its cost of equity using the bond-yield-plus-risk-premium approach. What is the firm's cost of equity using each of these three approaches? Do not round intermediate calculations. Round your answers to two decimal places. CAPM cost of equity: % Bond-Yield-Plus-Risk-Premium: % DCF cost of equity: % If you are equally confident of all three methods, then what is the best estimate of the firm’s cost of equity? Highest, Lowest, Or average.
Answer #1

a) Cost of Equity = Dividend at year end/Price + Growth =2.50/22+3.7%= 15.0636% or 15.06%

b) Cost of equity using CAPM = Risk Free Rate + beta *(Market Return - Risk Free Rate) = 4.7%+1.2*6% = 11.90%

c)Cost of equity using the bond-yield-plus-risk-premium-approach=9.48%+4% =13.48%

d) If you are equally confident of all three methods then Cost of common equity =(15.0636%+ 11.90%+13.48%)/3 =13.48%